OpenStudy (anonymous):

Find the coefficient of x^-6 in the expansion of (2x-3/x^2)^12

OpenStudy (mathmale):

Are you sure you mean \[x^{-6}?\]

OpenStudy (mathmale):

... and not \[x^6?\]

zepdrix (zepdrix):

Can you clarify the expression you wrote bri, is it this\[\large\rm \left(\frac{2x-3}{x^2}\right)^{12}\] or this\[\large\rm \left(2x-\frac{3}{x^2}\right)^{12}\]

OpenStudy (anonymous):

the latter

OpenStudy (anonymous):

ill be back in ten minutes, is that ok? ill just tag you when i come back real quick

zepdrix (zepdrix):

Ya do whatever you gotta do :d I might post some steps in the mean time.

zepdrix (zepdrix):

Binomial Theorem tells us we can write it this way,\[\large\rm (2x-3x^{-2})^{12}\quad=\sum_{k=0}^{12}\left(\begin{matrix}12 \\ k\end{matrix}\right)(2x)^{12-k}(-3x^{-2})^{k}\]We'll have to combine these x's I suppose, so it will be easier to determine which value of k corresponds to the x^{-6} term.

OpenStudy (anonymous):

ok I'm back! Yea, I kind of understand that. We have to find the general term, right?

zepdrix (zepdrix):

So if you split up the garbage you get something like this,\[\large\rm =\sum_{k=0}^{12}\left(\begin{matrix}12 \\ k\end{matrix}\right)2^{12-k}(-3)^k~~ x^{12-k}x^{-2k}\]And then we can apply exponent rule to combine the x's, ya?

OpenStudy (anonymous):

kind of..lemme just look at it for a sec to understand it all lol

zepdrix (zepdrix):

Ya lemme slow down a sec, as a first step I'm applying Binomial Theorem,\[\large\rm (a+b)^n\quad=\sum_{k=0}^n \left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k}b^k\]

OpenStudy (anonymous):

OHH ok haha I get it. the thing is, I'm in IB and SL math uses r instead of k

zepdrix (zepdrix):

Ah :)

OpenStudy (anonymous):

but ye ok, i get it. also, thanks for separating it all, it makes waaaay more sense that way

zepdrix (zepdrix):

R's are better? :)\[\large\rm =\sum_{r=0}^{12}\left(\begin{matrix}12 \\ r\end{matrix}\right)2^{12-r}(-3)^r~~ x^{12-r}x^{-2r}\]So combining our x's gives us,\[\large\rm =\sum_{r=0}^{12}\left(\begin{matrix}12 \\ r\end{matrix}\right)2^{12-r}(-3)^r~~ x^{12-3r}\]

zepdrix (zepdrix):

Which value of r gives you the x^{-6} term?\[\large\rm (stuff)x^{12-3r}=(stuff)x^{-6}\]

OpenStudy (anonymous):

umm..it's the 12-3r right?

OpenStudy (anonymous):

cos like thats the general term right..

zepdrix (zepdrix):

You have to figure out which numerical value of r gives you the x^{-6} on the left side of that equation.

OpenStudy (anonymous):

oh ok lol I'm totally lost then

zepdrix (zepdrix):

You're essentially boiling it down to this: \(\large\rm 12-3r=-6\) and solving for r. You simply want to know what value of r gives you that -6 for the exponent.

OpenStudy (anonymous):

ohh ok so then it's 6?

zepdrix (zepdrix):

Mmm good, that sounds right! :) So your answer will be the general term,\[\large\rm \left(\begin{matrix}12 \\ r\end{matrix}\right)2^{12-r}(-3)^r~~ x^{12-3r}\]evaluated at r=6

OpenStudy (anonymous):

waait wait wait the back of the book says the answer is something else lemme do the equation version, one sec

OpenStudy (anonymous):

\[\left(\begin{matrix}12 \\ 6\end{matrix}\right) * 2^{6}* (-3)^6\]

OpenStudy (anonymous):

omg wait

OpenStudy (anonymous):

no i get it lol

zepdrix (zepdrix):

Oh they only wanted the coefficient :) my bad. So we should have left off the x^{-6}.

zepdrix (zepdrix):

The rest matches up though, yes? :D

OpenStudy (anonymous):

yeah it makes sense! thank you soo much. would you mind helping me with one more? i could make another question if you want so you could get 2 best respoinses

zepdrix (zepdrix):

ya i can try c:

OpenStudy (anonymous):

thank u haha ill make the new q now